# Poincaré series of semigroups

### v1

@@ 4. Alexander duals
Consider a finite set $E$ and a simplicial subcomplex $T$ of the simplex $\mathcal{P}(E)$ with set of vertices $E$. The subset
@@
are called the Alexander dual and the relative Alexander dual of $T$. The properties of the Alexander duality we will need are summarized in the following result.
@@
$\overline{T}^{\vee}=\{K\cup(E\backslash E^{\prime})\,|\,K\in T^{\vee}\}\cup\{C\cup F\,|\,C\subset E^{\prime},\,F\subset E\backslash E^{\prime},\ F\neq E\backslash E^{\prime}\}$
@@
Statement (4) considers the special case in which $T$ consists only of the empty face, in that case $T^{\vee}$ is not defined as $E^{\prime}$ is empty too. However, $\overline{T}^{\vee}$ is defined and it is a (#$\#E-2$)-topological sphere, so $\tilde{\chi}(\overline{T}^{\vee})=(-1)^{\#E-2}=(-1)^{\#E}$. Then one needs to define $\tilde{\chi}(T^{\vee})=1$ in order that (3) be again true for this case.
@@
Let $S$ be a finitely generated positive cancelative commutative semigroup and $E\subset S$ a finite subset. Then one has the following two formulas for its Poincaré series $P=P_{S}\in\hat{R}.$
@@
First, #$\#E^{\prime}=0$ only occur for $m\in Q$. The dual $T_{m}^{\vee}$ is not defined, but the value $\tilde{\chi}(T_{m}^{\vee})$ is defined and equal to 1. If #$\#E^{\prime}=1$, $E^{\prime}=\{e\}$ then $T_{m}$ is one point and $T_{m}^{\vee}=\emptyset$. If #$\#E^{\prime}=2$, $T_{m}$ is either two isolated points or one segment, in the first case $T_{m}^{\vee}=\{\emptyset\}$, in the second $T_{m}^{\vee}=\emptyset$. If #$\#E^{\prime}=3$, the five topological types for $T_{m}$ listed in section $1$ have respective duals $T_{m}^{\vee}$ listed in figure 3.
@@
5. Symmetry
[…]
The notion of symmetric semigroup for higher dimensions is not so much developed in the literature, for instance see [GMSV] where a notion of symmetric semigroup is considered for toric $S$ having a finite complement in $G(S)\cap C(S)$. The difficulty for dimension bigger than 1 is that a good notion of gaps, is still missing.
[…]
Consider the Poincaré series $P\in\mathbb{Z}[[t]]$ of $S$, choose a non empty finite subset $E\in S\backslash\{0\}$ and for it also consider the polynomial $p=\prod_{e\in E}(1-t^{e})P$, the simplicial complexes $T_{m}$, the Apery set $Q$ and the semigroup element $g=f+e_{E}$. Two elements $m,m^{\prime}\in G(S)=\mathbb{Z}$ such that $m+m^{\prime}=g$ are said to be symmetric one to each other. Denote also by $Q_{E}$ the Apery set with respect to the choice $\{e_{E}\}$. Also denote
$\overline{P}$ $=P(1/t)\in\mathbb{Z}[[t,t^{-1}]]$
$\overline{p}$ $=p(1/t)\in\mathbb{Z}[t,t^{-1}]$
and by
$P_{G(S)}=\sum_{i\in\mathbb{Z}}t^{i}$
the Poincaré series of the group $G(S)=\mathbb{Z}$.
Proposition 5.1.
For a numerical semigroup $S$ the following conditions are equivalent:
  1. (1)
    $S$ is symmetric
  2. (2)
    $P+t^{f}\overline{P}=P_{G(S)}$
  3. (3)
    $p+(-1)^{\#E}t^{g}\overline{p}=0$
  4. (4)
    If $m,m^{\prime}$ are symmetric then $\overline{T}_{m}^{\vee}=T_{m^{\prime}}$
  5. (5)
    If $m,m^{\prime}$ are symmetric then $m\in Q_{E}$ iff $m^{\prime}\in Q_{E}$.
Proof.
$P$ is the sum expression of $t^{m}$ that $m\in S$. $t^{f}\overline{P}$ the sum expression of $t^{f-n}$ that $n\in S$. If $S$ is symmetric, above second expression is that of the $t^{m}$ such that $\𝑆m\in\mathbb{Z}\backslash S$. So $P+t^{f}\overline{P}$ is the sum expression of all $t^{m}$, i.e. $P_{G(S)}$. This shows (1$\Rightarrow$ (2).
Now multiplying times $q=\prod_{e\in E}(1-t^{e})\in R$ both sides of the equality in (2), and taking into account that $1-t^{e}=-t^{e}(1-t^{-e})$, that the right side becomes 0 and that $g=f+e_{E}$, one shows (2$\Rightarrow$(3). Moreover, multiplying times $q^{-1}\in\hat{R}$ both sides of the equality in (3) one gets the following equality in $\hat{R}$:
$\sum_{m\in S}t^{m}+\sum_{\begin{subarray}{c}n\in S\\ n<f\end{subarray}}t^{f-n}-(1+t+t^{2}+\cdots)=0.$
This proves that $f-m\notin S$ iff $m\in S$ therefore (3$\Rightarrow$(1).
Assume again (a), and $m,m^{\prime}$ with $m+m^{\prime}=g$. Then
$K\in\overline{T}_{m}^{\vee}\Leftrightarrow m-e_{E\backslash K}\notin S\Leftrightarrow f-m+e_{E\backslash K}=m^{\prime}-e_{K}\in S\Leftrightarrow K\in T_{m^{\prime}}$
hence (1$\Rightarrow$(4). If (4) is true, one has the equivalences below for symmetric elements $m,m^{\prime}$
$m^{\prime}\in Q_{E}\Leftrightarrow m^{\prime}\in S,\,m^{\prime}-e_{E}\notin S\Leftrightarrow\emptyset\in T_{m^{\prime}},\,E\notin T_{m^{\prime}}\Leftrightarrow\emptyset\in\overline{T}_{m}^{\vee},\,E\notin\overline{T}_{m}^{\vee}\Leftrightarrow$
$m-e_{E}\notin S,\,m\in S\Leftrightarrow m\in Q_{E^{\prime}}$
so (4$\Rightarrow$(5).
Finally, let us prove (5$\Rightarrow$(1). For it, take $m\notin S$. Since $S$ has a conductor, there exists $\ell\geq 1$ such that $m+\ell e_{E}\in S$ and $m+(\ell+1)e_{E}\notin S$, in particular, $m+\ell e_{E}\in Q_{E}$. By (5), one has $g-m-\ell e_{E}=f-m-(\ell-1)e_{E}\in Q_{E}$ and hence $f-m\in S$. Thus, $S$ is symmetric as required. ∎
Above $m,m^{\prime}$ are not excluded to be elements of $G(S)$ not in $S$, assuming that their simplicial complex is $\emptyset$, i.e. the complex with no faces. In fact, $\emptyset$ is relevant in Alexander duality as it is the dual of the full simplex.
Looking to the different statements of the proposition, one can think that they could be concrete tentatives to give respective notions of symmetry for semigroups of dimension bigger than 1.
However (1) and (2) are not easy to explore by geometrical reasons as $G(S)\backslash S$ is quite big. On the other hand (5) is an empty condition in dimension $d>1$ as $Q_{E}$ is an infinite set and, from the statement, one deduces that $Q_{E}$ is finite as $\lambda(m)\leq\lambda(g)$ for all $m\in Q_{E}$ and the fibers of $\lambda$ are finite, $\lambda$ be as in section 1. Finally (4) is equivalent to (5), so it is an empty statement for $d>1$ too.
Thus, it is the condition in statement (3) that is usually considered as symmetry, in the quite general situation of graded rings in commutative algebra, where it corresponds to the condition on the rings to be Gorenstein, as detailed below in the coming section.
For numerical semigroups it is well known that $S$ is symmetric iff $R\otimes_{\mathbb{Z}}\mathbb{K}=\mathbb{K}[S]$ is Gorenstein for any field $\mathbb{K}$, see [K].
The same happens for semigroups of reduced curve singularities which are not finitely generated, for which a good notion of symmetry is established.
@@ 6. Syzygies
We exclude $j=\#E-1$, i.e. $T=\mathcal{P}(E)$ since its homology vanishes.
@@
On the other hand, the homology type of every simplicial complex $T$ can be realized as that of one $T_{m}$ for some concrete semigroup $S$ and $m\in S$, as shown in [BH]. To see it, take the subsemigroup $S$ of #$\mathbb{Z}^{\#E+1}$ generated by the elements of type $(0,...,0,1,0,...,0)$ where the last coordinate is 0 and the integer 1 is the coordinate of the label of a vertex of $T$ for a chosen labeling, and by elements of type $(0,1,...,1,0,...,1)$ where 1 is the last coordinate and the other coordinates 1 are those of the vertices of a maximal face of $T$. Then, one can easily check that the element $m=(1,...,1)\in\mathbb{Z}^{\#E+1}$ belongs to $S$ and $T_{m}$ has the homotopy type of $T$.
@@
where $b_{j},\,-1\leq j\leq s$ are the Betti numbers $b_{j}>0$. Its length $s$ is called the homological dimension and, according to the Auslander-Buchbaum theorem, its complement $r=\#E-s$ is nothing but the depth of $\mathbb{K}[S]$. For simplicity, we will assume that $E$ generates all the semigroup $S$, i.e. $Q=\{0\}$, $b_{-1}=1$.
@@
The graded Betti numbers are directly related to our topological tools by the equalities
@@
      The algebra $\mathbb{K}[S]$ is Gorenstein.
@@
      One has $b_{j,m}=b_{s-j,m^{\prime}}$ for $m,m^{\prime}$ symmetric with respect to some $g\in S$. In particular $b_{-1}=b_{s,g}=1$.
@@
      For some $g\in S$, $p$ satisfies the equation $p+(-1)^{\#E+d-1}t^{g}\overline{p}=0$
@@
For general choices of $E$ generating the cone $C(S)$, the first term of (1) is #$\#Q$. We have assumed that $E$ generates $S$, thus $Q=\{0\}$, $N_{Q}=1$. Statements (1), (2b), (2c) for a choice $E$ generating the cone $C(S)$ can be deduced from the corresponding to a larger choice $E\cup A$ which generates the semigroup $S$.
If an element in conditions (2b) and (2c) exists, then it is unique and the same for both.
Condition (2c) extends, for $d>1$, condition (c) in proposition of above section. In fact, one can easily see that $f=g-e_{E}$ does not depend on $E$ and one has
@@
which can be seen as another equivalent condition to those in (2).
The semigroups of complete intersection are Gorenstein. For the choice of $E$ generating the semigroup $S$ one has $Q=\{0\}$, and the kernel of the first map of the resolution $R_{E}\rightarrow\mathbb{K}[S]$ is generated by $b_{0}\geq\#E-d$ $S$-homogeneous polynomials of degrees $m$ for which $b_{0,m}\neq 0$, properly counted. When $b_{0}=\#E-d$, $\mathbb{K}[S]$ is said to be a complete intersection. In that case, denote by $C$ the set of those degrees, each $m$ repeated $b_{0,m}$ times, and $g=\sum_{c\in C}c$.
@@
Finally, we will show some relations of the depth $r$ and the Poincaré series. Assume again here that $E$ generates the cone $C(S)$. In [CG] Theorem 4.1, it is proved that $r$ is the integer satisfying the following two conditions
$\tilde{h}_{\#E-r}(T_{m})=0\quad for\,all\quad m\in S$
$\tilde{h}_{\#E-r-1}(T_{m})\neq 0\quad for\,some\quad m\in S.$
@@
$\tilde{h}_{r-3}(\overline{T}_{m}^{\vee})=0\quad for\,all\quad m\in S$
$\tilde{h}_{r-2}(\overline{T}_{m}^{\vee})\neq 0\quad for\,some\quad m\in S$
@@
Above, for sphere, we mean a (#$(\#E-2)$-dimensional sphere. Also the dual complexes can be visualized in terms of the Apery set $Q_{E}$ of the single choice $\{e_{E}\}$. In fact, one has $J\in\overline{T}_{m}^{\vee}$ iff $m+e_{J}\in Q_{E}$.
Next, we will relate the depth $r$ with the key combinatorial sets of the semigroup.
@@
$D^{J}=\emptyset\quad for\,all\quad J\subset E\quad\#J\geq\#E-r^{\prime}+2$
$D^{J}\neq\emptyset\quad for\,some\quad J\subset E\quad\#J=\#E-r^{\prime}+1.$
@@
To prove (3), take $J\subset E$, #$\#J=\#E-r^{\prime}+1$, such that $D^{J}\neq\emptyset$, and an element $m\in D^{J}$. One has $J\subset supp(m)$ and $m-e_{J}\notin S$. If $J$ is a (#$(\#E-r^{\prime}-1)$-sphere of $T_{m}$, the result follows from (2). Otherwise, there exists $e\in supp(m)$ such that $m-e_{J}+e=m+e-e_{J}\notin S$. Now continue with $m+e$ instead $m$. If $J$ is a (#$(\#E-r^{\prime}-1)$-sphere of $T_{m+e}$ the result follows from (2), if not repeat the argument. This procedure generates a increasing sequence of complexes $T_{n}$ with $J\notin T_{n}$ and $J\subset supp(n)$. Since $D$ is finite, $J$ needs to be a (#$(\#E-r^{\prime}-1)$-sphere of $T_{n}$ for some $T_{n}$ in the sequence. This completes the proof of the theorem applying again (2). ∎
@@
The main results of [CG] show that also $P^{h}$ is determined, not only by the topological tools, but also by the combinatorial ones. Thus, we have the following result.

### v2

@@ 4. Alexander duals
Consider a nonempty finite set $E$ and a simplicial subcomplex $T$ of the simplex $\mathcal{P}(E)$ consisting of all subsets (parts) of $E$ with set of vertices $E$. The subset
@@
are called the Alexander dual and the relative Alexander dual of $T$. The properties of the Alexander duality we will need are summarized in the following result (see also [BM]).
@@
$\overline{T}^{\vee}=\{K\cup(E\backslash E^{\prime})\,|\,K\in T^{\vee}\}\cup\{C\cup F\,|\,C\subset E^{\prime},\,F\subset E\backslash E^{\prime},\ F\neq E\backslash E^{\prime}\}.$
@@
Statement (4) considers the special case in which $T$ consists only of the empty face, in that case $T^{\vee}$ is not defined as $E^{\prime}$ is empty too. However, $\overline{T}^{\vee}$ is defined and it is a (#$\#E-2$)-topological sphere, so $\tilde{\chi}(\overline{T}^{\vee})=(-1)^{\#E-2}=(-1)^{\#E}$. Then one needs to define $\tilde{\chi}(T^{\vee})=1$ in order for (3) to be true again.
@@
Let $S$ be a finitely generated positive cancelative commutative semigroup and $E\subset S\backslash\{0\}$ a nonempty finite subset. Then one has the following two formulas for its Poincaré series $P=P_{S}\in\hat{R}.$
@@
First, #$\#E^{\prime}=0$ only occur for $m\in Q$. The dual $T_{m}^{\vee}$ is not defined, but, as said before, the value $\tilde{\chi}(T_{m}^{\vee})$ is defined and equal to 1. If #$\#E^{\prime}=1$, $E^{\prime}=\{e\}$ then $T_{m}$ is one point and $T_{m}^{\vee}=\emptyset$. If #$\#E^{\prime}=2$, $T_{m}$ is either two isolated points or one segment, in the first case $T_{m}^{\vee}=\{\emptyset\}$, in the second $T_{m}^{\vee}=\emptyset$. If #$\#E^{\prime}=3$, the five types for $T_{m}$ are listed among those in figure 2 and have respective duals $T_{m}^{\vee}$ listed in figure 3.
@@ 5. Syzygies
We exclude $j=\#E-1$, i.e., $T=\mathcal{P}(E)$ since its homology vanishes.
@@
On the other hand, the homology type of every simplicial complex $T$ with $E^{\prime}=E$ can be realized as that of one $T_{m}$ for some concrete semigroup $S$ and $m\in S$, as shown in [BH]. To see it, take the subsemigroup $S$ of #$\mathbb{Z}^{\#E+1}$ generated by the elements of type $(0,...,0,1,0,...,0)$ where the last coordinate is 0 and the integer 1 is the coordinate of the label of a vertex of $T$ for a chosen labeling, and by elements of type $(0,1,...,1,0,...,1)$ where 1 is the last coordinate and the other coordinates 1 are those of the vertices of a maximal face of $T$. Then, one can easily check that the element $m=(1,...,1)\in\mathbb{Z}^{\#E+1}$ belongs to $S$ and $T_{m}$ has the homotopy type of $T$.
@@
where $b_{j},\,-1\leq j\leq s$ are the Betti numbers $b_{j}>0$. Its length $s$ is called the homological dimension and, according to the Auslander-Buchbaum Theorem, its complement $r=\#E-s$ is nothing but the depth of $\mathbb{K}[S]$. For simplicity, we will assume that $E$ generates all the semigroup $S$, i.e., $Q=\{0\}$, $b_{-1}=1$.
@@
The graded Betti numbers are directly related to our combinatorial tools by the equalities
@@
      The algebra $\mathbb{K}[S]$ is Gorenstein
@@
      One has $b_{j,m}=b_{s-j,m^{\prime}}$ for all pairs $m,m^{\prime}$ symmetric with respect to some concrete $g\in S$. In particular $b_{-1}=b_{s,g}=1$
@@
      For some $g\in S$, $p$ satisfies the equation $p+(-1)^{\#E+d-1}t^{g}\overline{p}=0$ where $\overline{p}(t)=p(1/t)$.
@@
For general choices of $E$ generating the cone $C(S)$, the independent term of the polinomial in (1) is #$\#Q$. We have assumed that $E$ generates $S$, thus $Q=\{0\}$, $N_{Q}=1$. Statements (1), (2b), (2c) for a choice $E$ generating the cone $C(S)$ can be deduced from the corresponding to a larger choice $E\cup A$ which generates the semigroup $S$.
If an element $g$ in conditions (2b) and (2c) exists, then it is unique and the same for both.
Condition (2c) extends, for $d>1$, the usual condition for the symmetry of numerical semigroups. In fact, in that case, one can easily see that $f=g-e_{E}$ does not depend on $E$ and one has
@@
where $\overline{P}(t)=P(1/t)$, $c=f+1$ is the conductor of $S$.
The semigroups of complete intersections are Gorenstein. For the choice of $E$ generating the semigroup $S$ one has $Q=\{0\}$, and the kernel of the first map of the resolution $R_{E}\rightarrow\mathbb{K}[S]$ is generated by $b_{0}\geq\#E-d$ $S$-homogeneous polynomials of degrees $m$ for which $b_{0,m}\neq 0$, properly counted. When $b_{0}=\#E-d$, $\mathbb{K}[S]$ is said to be a complete intersection. In that case, denote by $C$ the set of those degrees, each $m$ repeated $b_{0,m}$ times, and $g=\sum_{c\in C}c$.
@@
Finally, we will show some relations of the depth $r$ and the Poincaré series. Assume again here that $E$ generates the cone $C(S)$. In [CG, Theorem 4.1], it is proved that $r$ is the integer satisfying the following two conditions
$\tilde{h}_{\#E-r}(T_{m})=0\text{ for all }m\in S$
$\tilde{h}_{\#E-r-1}(T_{m})\neq 0\text{ for some }m\in S.$
@@
$\tilde{h}_{r-3}(\overline{T}_{m}^{\vee})=0\text{ for all }m\in S$
$\tilde{h}_{r-2}(\overline{T}_{m}^{\vee})\neq 0\text{ for some }m\in S$
@@
Above, by sphere, we mean a (#$(\#E-2)$-dimensional sphere. Also the dual complexes can be visualized in terms of the Apéry set $Q_{E}$ of the single choice $\{e_{E}\}$. In fact, one has $J\in\overline{T}_{m}^{\vee}$ iff $m+e_{J}\in Q_{E}$.
Above results are reformulations of known results in terms of the Alexander duals of the combinatorial tools. Bellow, in theorem 5.3 we will give new results in terms of the set theoretical tools of our semigroups.
Next, we will relate the depth $r$ with the key set theoretical tools sets of the semigroup.
@@
$D^{J}=\emptyset\text{ for all }J\subset E\quad\#J\geq\#E-r^{\prime}+2$
$D^{J}\neq\emptyset\text{ for some }J\subset E\quad\#J=\#E-r^{\prime}+1.$
@@
To prove (3), take $J\subset E$, #$\#J=\#E-r^{\prime}+1$, such that $D^{J}\neq\emptyset$, and an element $m\in D^{J}$. One has $J\subset supp(m)$ and $m-e_{J}\notin S$. If $J$ is a (#$(\#E-r^{\prime}-1)$-sphere of $T_{m}$, the result follows from (2). Otherwise, there exists $e\in supp(m)$ such that $m-e_{J}+e=m+e-e_{J}\notin S$. Now continue with $m+e$ instead of $m$. If $J$ is a (#$(\#E-r^{\prime}-1)$-sphere of $T_{m+e}$ the result follows from (2), if not repeat the argument. This procedure generates a increasing sequence of complexes $T_{n}$ with $J\notin T_{n}$ and $J\subset supp(n)$. Since $D$ is finite, $J$ needs to be a (#$(\#E-r^{\prime}-1)$-sphere of $T_{n}$ for some $T_{n}$ in the sequence. This completes the proof of the theorem applying again (2). ∎
@@
The main results of [CG] show that also $P^{h}$ is determined, not only by the combinatorial tools, but also by the set theoretical ones. Thus, we have the following result.